# Explaining mountain pine beetle dynamics: From life history traits to large, episodic outbreaks

### v1

@@ Appendix D Supplementary Figures
Figure D.3: Sample time series of MPB-infested trees from British Columbia’s Aerial Overview Surveys. Each panel is a randomly selected $\text{1 km}^{2}$ patch from a subset of locations where cumulative infestations were in the upper quartile. To estimate the number of infested trees, we multiply the proportion of infested stems by Live aboveground biomass (units: $\frac{\text{Live aboveground biomass (units: Mg/ha)}}{10}$, which is an approximation for pine stem density that was validated by Koch et al., (2021, Appendix A), using the data and analyses of Nigh et al., (2008). The biomass data is sourced from Beaudoin et al., (2014).
Figure D.4: In 1970, emergence density was high but subsequent attack density was low. Anomalous weather conditions during the flight period could be responsible for this pattern, but the relevant weather variables were unexceptional in 1970 (the bold, teal line). The maximum summer temperature was consistently greater than 18.3°C, the lower limit for beetle flight (McCambridge,, 1971). The estimated wind speed often exceeded 8 km/hour, the proposed upper threshold for beetle flight (Gray et al.,, 1972); however, the wind speeds in 1970 were not anomalous in comparison to the other outbreak years. Weather data was obtained from the Biosim program (Régnière et al.,, 2014).
Figure D.5: In 1970, emergence density was high but subsequent attack density was low. The decline could theoretically be attributed to poor environmental conditions that weakened the beetles, despite not being lethal (because they still emerged). However, the developmental period leading up to 1970 had unexceptional weather conditions. The minimum winter temperature is related to larval mortality (here, sub-lethal damage), and degree days are related to emergence and univoltinism. The minimum winter temperature is approximately -30°C, which is substantially greater than the proposed threshold of -40°C (Safranyik,, 1998) for significant overwintering mortality. There are approximately 1250 degree days from Aug. 1 to Jul. 31, far beyond the proposed threshold of 833 for the population to be univoltine (Reid,, 1962). Weather data was obtained from the Biosim program (Régnière et al.,, 2014).
@@
Figure D.6: The stable epidemic equilibrium (i.e., the intersection of the nullclines with the highest beetle density) has a very small basin of attraction. All other initial conditions lead to the endemic equilibrium.

### v2

@@ Appendix E Additional supplementary figures
Figure E.1: Sample time series of MPB-infested trees from British Columbia’s Aerial Overview Surveys. Each panel is a randomly selected $\text{1 km}^{2}$ patch from a subset of locations where cumulative infestations were in the upper quartile. To estimate the number of infested trees, we multiply the proportion of infested stems by Live aboveground biomass (units: $\frac{\text{Live aboveground biomass (units: Mg/ha)}}{10}$, which is an approximation for pine stem density that was validated by Koch et al., (2021, Appendix A), using the data and analyses of Nigh et al., (2008). The biomass data is sourced from Beaudoin et al., (2014).
Figure E.2: Graphical representation of our inferential approach. First, we develop and validate a realistic model of MPB dynamics. Second, we perform a causal analysis in which we turn biological features on or off. In the example of the causal analysis, the red crosses respectively indicate that feature # 1 has been off, and that the simulated system does not match reality with respect to Metric #2 (e.g., outbreak periodicity). The dashed red arrow indicates that feature #1 is responsible for realistic values of Metric #2.
Figure E.3: Population maps provide evidence for compensatory density dependence at the tree level. The data comes from Alberta’s disk surveys (described by Government of Alberta,, 2016, and published by Goodsman et al.,, 2018). The estimate for attack density at time $t+1$ involves the constants $(2/3)$ (the average proportion of the brood that will be female) and $0.3$ (an estimate of dispersal-phase mortality from the Klein et al., (1978) dataset, see Fig. E.4).
Figure E.4: An approximation of dispersal mortality as a function of time in the Klein et al., data. The dispersal mortality is approximated as the quotient of gallery starts and preceding emergence holes. The quotient captures both dispersal mortality and mortality from unsuccessful attacks (i.e., the Allee effect, Eq. 3). Therefore, the approximation will be better when emergence density is larger (i.e., before 1971), because the Allee effect will be less important and the relative influence of inter-patch dispersal will be less important. The first three years can be used to approximate the baseline dispersal mortality (i.e., $1-s_{0}$ in Eq. 1).
Figure E.5: Frequency distributions of attack density. The Model posterior category gives the marginal posterior of $1/\gamma$ from our statistical model of the Klein et al., outbreak. The AB disk data category contains the within-site means of attack density from 300 different sites across Alberta (Government of Alberta,, 2016). The Other studies category includes estimates obtained via entry holes or gallery starts from a number of studies, mostly United States Forest Service report: Tishmack et al., (2005); Mccambridge, (1967); Negrón, (2018); Reid, (1963); Raffa and Berryman, (1983); Rasmussen, (1980); Shepard, (1965); Knight, (1959); De Leon, (1939); Safranyik, (1988); Safranyik and Vithayasai, (1971); Safranyik, (1968); Parker, (1979); Schmid, (1972); Blackman, (1931); Whiteside, (1937); Washburn and Cole, (1959); author N/A, (1963); Peterman, (1974); Klein et al., (1978); Safranyik and Linton, (1991).
Figure E.6: Frequency distributions of female emergence density. Females are estimated with (2/3)×(3/2)×(2/3$\times$(3/2$\times$( 2 / 3 ) × ( 3 / 2 ) × emergence holes, or with (2/3)×(2/3$\times$( 2 / 3 ) × emerging beetles or spring brood size. Note that (2/3) is the typical proportion of females in a brood (Reid,, 1962; Cole et al.,, 1976), and (3/2) is the average number of beetles emerging from each hole (Safranyik and Linton,, 1985; Peterman,, 1974, Fig. 13). The Model posterior category gives the marginal posterior of $\zeta_{0}+\zeta_{1}*10$, the emergence density on a 10-inch DBH tree; 10 inches is approximately the average DBH of MPB-killed trees in the Klein et al., dataset. The AB disk data category contains estimates of emergence from 300 different sites across Alberta (Government of Alberta,, 2016). The Other studies category includes estimates from a number of studies, mostly United States Forest Service reports: Tishmack et al., (2005); Negrón, (2018); Reid, (1963); Raffa and Berryman, (1983); Rasmussen, (1980); De Leon, (1939); Safranyik, (1988); Cole, (1969); Schmid, (1972); Whiteside, (1937); Beal, (1938); Peterman, (1974); Klein et al., (1978); Safranyik and Linton, (1985, 1991).
@@
Figure E.7: The stable epidemic equilibrium (i.e., the intersection of the nullclines with the highest beetle density) has a very small basin of attraction. All other initial conditions lead to the endemic equilibrium.